OpenStudy (anonymous):

2,000(A - 5) = 110,000; A equals

OpenStudy (anonymous):

solving for a is your goal divide both sides by 2000 add 5 to both sides

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Wait Juanita aren't you supposed to distribute the 2000 to A and - 5?

OpenStudy (anonymous):

dont have to...since the parenthesis is there you can divide to move w/o distrubuting

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

use distrubutive property to get 2000a -10000 = 11000 then add then divide by 2000 to get

OpenStudy (anonymous):

a-5 = 55 a-5+5 = 55+5 a = 60

OpenStudy (anonymous):

both ways get same answer

OpenStudy (anonymous):

sure does - i just try to work with smaller numbers as much as i can. thats why i said "you dont have to"